OpenStudy (anonymous):

there are approximately 3lbs of muscle for every 7 lbs of body weight of a certain animal. for a 161lb animal how much of the weight is muscle

OpenStudy (anonymous):

3lbs muscle/7lbs weight = x muscle / 161 weight 3/7 = x/161 3*161=7x 483=7x x=69 so 69 lbs of muscle

OpenStudy (anonymous):

checked it a couple times and i have to agree with r3dd.